Circular No. 6921

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998ca, 1998ci, 1998cj N. Regnault and O. Perdereau, Laboratoire de l’Accélérateur Linéaire,

Université de Paris-Sud, on behalf of the EROS collaboration (cf. IAUC 6762, 6904), report the discovery of two apparent supernovae in the course of the automated supernova search with the 1-m Marly telescope (+ CCD) at La Silla. SN 1998ci, discovered at V ∼ 18 on images taken on May 19, is located at α = 12^{h}41^{m}01^{s}.12, δ = −6^{o}39′48′′.2 (equinox 2000.0), which is very close to the center of the anonymous host galaxy. SN 1998ci was not present on a frame taken on 1997 Apr. 4, but it was confirmed on frames taken on 1998 May 20.

SN 1998cj, discovered on images taken on May 21 at V ∼ 17.2, is located at α = 12^{h}19^{m}43^{s}.30, δ = −12^{o}13′41′′.8 (equinox 2000.0), which is 9′′.4 east and 10′′.8 south of the center of NGC 4263. This star was not present on a CCD frame taken on 1997 Feb. 7. SN 1998cj was confirmed on CCD frames taken on 1998 May 26.

G. Richards, University of Chicago; A. Kim, Physique Corpusculaire et Cosmologie, Collège de France; P. Nugent, Lawrence Berkeley Laboratory; H. Newberg and S. Kent, Fermi National Accelerator Laboratory, and the EROS collaboration report that spectra were obtained on May 27 UT with the 3.5-m ARC Telescope (+ Double Imaging Spectrograph) at Apache Point Observatory. SN 1998cj is of type II, more than a month past maximum light. SN 1998ci is a significantly reddened type-Ia supernova (like SN 1991T) at z = 0.025. SN 1998ca (IAUC 6904) is a post-maximum type-II event, within a month of maximum light, at z = 0.028. The redshifts (± 0.001) were estimated from absorption features in the galaxy spectra.

SUPERNOVA 1998bu IN NGC 3368 R. Evans and W. Wild, University of Chicago, report additional mag-

nitudes (cf. IAUC 6902) for SN 1998bu, obtained with the Yerkes 1-m telescope (estimated error ± 0.15 mag): May 21.10 UT, I = 12.0; 21.11, R = 11.8; 21.13, B = 12.4, V = 12.0. Absolute calibration was done using HR 1433, HR 4983, and HR 3427.

CCD photometry by D. Hanžl, Brno, Czech Republic (0.40-m reflector), based on comparison star GSC 849.733 (Tycho magnitudes V = 9.73 ± 0.03, B − V = +0.636 ± 0.04): May 10.835 UT, V = 12.83 ± 0.04 (revision of value on IAUC 6899); 11.861, V = 12.59 ± 0.02; 14.848, V = 12.25 ± 0.02, B − V = +0.40 ± 0.02; 20.857, V = 12.01 ± 0.04, B − V = +0.48 ± 0.04.

1998 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green